PH 142 Total: 25 Due: 02/01/19 at Question 3-6 points (a) Calculate the electric field in a cell membrane with a thickness of 8nm when the membrane potential is -70 mV. (2 points) (b) If a Sodium ion, Na, moves from inside the cell to outside the cell, calculate the change in the ions electric potential energy. Does it increase or decrease? (2 Points) (e) In one cycle the sodium-potassium pump removes 3 sodium ions from the cell while importing 2 potassium ions, K. What is the total change in electric potential energy for one cycle of the pump? (2 Points) Question 4-6 points lon channels that switch open and close as the potential difference across the cell membrane changes are an essential component of the cell membrane. Most such channels contain a molecular ion with a significant electrical charge at their core. Suppose a protein contains a molecular ion with a charge of-12e and that this protein is embedded in a cell membrane that has a thickness of 9 nm. What are the magnitude and direction of electric force on the ion under the following conditions: to
Answer #1

3) a) electric field = voltage / separation

E = V/d

= 0.07 / 8*10^-9

= 8.75*10^6 N/C

(b) it will decrease

EPE = qV = 1.6*10^-19*-0.07 = 1.12*10^-20 J